OpenStudy (fixxer):

Lets say I have a circle centred at X=-1 The radius is 1, how can I express r? The area integral will need to be expressed in terms of drdt where t ranges from Pi/2 to 3Pi/2

OpenStudy (fixxer):



OpenStudy (fixxer):

 Unit circle t: 0 to 2Pi but in the other case when r: 0 to -2 t:Pi/2 to 3Pi/2

OpenStudy (fixxer):

\[\int\limits \!r\,{\rm d}[r={0\ldots -2\,\cos \left( t \right) },t={\pi/2 \ldots 3/2\,\pi}]\] I know excplicit for this circle, but more in general, how can i express r if the circle isnt centred at orgin

OpenStudy (holsteremission):

Consult the standard form of the circle equation: \[(x-a)^2+(y-b)^2=r^2\]is a circle centered at \((a,b)\) with radius \(r\). Converting to polar coordinates with radius \(1\) yields \[\begin{align*} (x+1)^2+y^2=1\implies (r\cos\theta+1)^2+(r\sin\theta)&=1\\[1ex] r^2\cos^2\theta+2r\cos\theta+1+r^2\sin^2\theta&=1\\[1ex] r^2+2r\cos\theta&=0 \end{align*}\]We require \(r>0\), so we can divide through by \(r\) to get \[r+2\cos\theta=0\implies r=-2\cos\theta\]

OpenStudy (holsteremission):

And in general, with arbitrary radius \(R\) and center \((a,b)\), you end up with \[\begin{align*} R^2&=(r\cos\theta-a)^2+(r\sin\theta-b)^2\\[1ex] R^2&=r^2\cos^2\theta-2ar\cos\theta+a^2+r^2\sin^2\theta-2br\sin\theta+b^2\\[1ex] R^2-a^2-b^2&=r^2-2r(a\cos\theta+b\sin\theta)\\[1ex] R^2-a^2-b^2&=\bigg(r-(a\cos\theta+b\sin\theta)\bigg)^2-(a\cos\theta+b\sin\theta)^2 \end{align*}\]and so on.

OpenStudy (fixxer):

Amaizing, you never fail to impress :)